# Breaking the News: A LLM-based Game where Players Act as Influencer or Debunker for Raising Awareness About Misinformation

### v1

@@ 6.3.1. Balance Between Freedom and Guidance
In our game, we sought to challenge the existing choice-based approach in misinformation education games by providing users with free-form input. We found this approach triggers reflection, which helps to build hands-on experience and make the game more enjoyable. Yet, we also found that it relies on players’ existing knowledge of misinformation. For example, players might incorporate unverified information they’ve encountered on social media into the game, which is specifically problematic for the debunker role. While we provided the players with comprehensible instructions to guide their role’s actions (how to act as a debunker or an evil influencer), it would be better to incorporate more context-specific tips in each stage of the game to help users explore different ways of winning the game and deepening their learning. We suggest using the approach used in the free-input educational interventions (e.g. (Bryfczynski, 2012)) to build clear, understandable criteria for free-form answers. These will not stop creativity but help people tailor their answers to the context of the game. We also suggest adding a preliminary stage to the game in the form of a simple choice-based questionnaire, helping people understand the role and the guidelines for the role. For example, we can introduce to the individual playing the debunker role the Debunking Ethics Standards (EDMO, 2024; Agence France-Presse, 2023; European External Action Service, 2021), and ask them a series of questions about understanding the content of the guidelines before starting the main game. In addition, we can incorporate an additional AI-based mechanic to check the quality of answers, not only in terms of its influentiality, but also of its coherence with debunking practices guidelines. Similarly, we can demonstrate a ”Score of Manipulativeness” to the influencer role. This better shows us how the ideas players implement into the message can be judged, from the point of view of dealing with misinformation.
@@ 6.3.2. Replayability and Feedback
One of the critical challenges in serious games is maintaining replayability, as this is important to facilitate the learning process (Adetunji and Ade-Ibijola, 2024) and making interventions more sustainable (Silveira, 2016). Moreover, a lack of replayability in educational games can limit both educational and behavioral change (Epstein et al., 2021). To address this, we incorporated elements designed to ensure the game can be played multiple times, such as offering two distinct player roles and providing a free-form input mechanism that allows participants to explore a broader decision space. However, we recommend that future interventions consider additional methods to further enhance replayability.
One such approach involves leveraging opportunities to introduce various characters to represent public opinion. In our game, we observed that players finds characters’ feedback more engaging when this resonates with them personally. By introducing more characters, or by allowing players to customize characters to better reflect their own experience, the game could encourage players to return and interact with new characters. The result would be a more engaging experience. For instance, research has shown that debunking misinformation often occurs within families and can sometimes lead to conflict (Scott et al., 2023). In such a scenario, players could customize a character based on their previous experiences with family members, thereby practicing their own debunking strategies in a risk-free environment.
While in-game feedback guides player behavior during play, post-game debriefing sessions help consolidate learning and improve future performance. Research has found that debriefing is a crucial opportunity for players to process and integrate their learning experiences(Crookall, 2014; Leder et al., 2024; Barzilai and Stadtler, 2024). After gameplay, we suggest arranging debriefing sessions that allow players to review their strategies, assess their effectiveness, and receive constructive feedback, potentially improving learning outcomes. For instance, after a session focused on combating misinformation, a post-game review might present an ideal debunking response or a well-supported counterargument. Such structured reflection enables players to internalize lessons and increases the likelihood that they will re-enter the game with newly gained insights, thereby reinforcing both learning and replayability.
@@ 6.4. Limitations and Future Work
We acknowledge the following limitations. Our participants were mainly from an East Asian country, which can limit the generalizability of our findings. Previous studies have shown that cultural contexts influence individuals’ perceptions of misinformation and interventions (Noman et al., 2024). For example, study revealed that Mexican and Spanish users were more likely to trust -related misinformation compared to users from Ireland, the UK and the USA (Roozenbeek et al., 2020a). At the same time, non-western participants demonstrated a significantly higher willingness to challenge misinformation. The effect of proposed misinformation design interventions was significantly greater than that seen in analysis of the behaviors and attitudes in the UK population(Noman et al., 2024). Previous studies also suggest that certain populations may face greater challenges in being able to critically evaluate information. For example, a large-scale study observed that Asian individuals encounter more difficulties in assessing health information from social media compared to other populations (Chandrasekaran et al., 2024). In addition, individuals from an Asian background are more likely to incorporate social media information into their health-related decisions, potentially increasing their susceptibility to misinformation (Chandrasekaran et al., 2024). Therefore, future work should explore the cross-cultural applicability of our findings in different cultural contexts. However, as the cultural background of our participants can potentially make them more vulnerable to misinformation than other populations, we believe that our intervention is valuable as providing insight into the outlook, activity and the relevance of certain interventions for these populations.
@@
Our sample was relatively homogeneous in age. A recent meta-analysis of articles about different intervention approaches showed that neither age nor gender significantly impacts the effectiveness of media literacy interventions (Lu et al., 2024). However, previous work has suggested that media literacy interventions designed for certain age groups (e.g., older adults and adolescents) achieved greater effects(Moore and Hancock, 2022; Hartwig et al., 2024). Future work should determine if our approach is efficient in other age groups of users and, if necessary, tailor scenarios to suit the various needs of different age groups.
@@
Thirdly, our study provided only a one-time intervention and observed immediate learning effects; Previous work showed that even a one-time interaction with an educational game can provide long-term improvement in misinformation recognition. For instance, Maertens et al. tested the game ”Bad News” and found that inoculation effects lasted for at least 13 weeks. This suggests the potential for the long-term effectiveness of active inoculation interventions with regular assessment(Maertens et al., 2021). Still, future research should include multiple time points to assess the long-term effectiveness of our game intervention. There should also be comparisons between one-time and multiple play sessions, with explorations of the impacts of players assuming different roles within the games. In this study, each player was limited to a single role, either a misinformation creator or a debunker. This resulted in different learning experiences depending on their assigned role. The primary reason for not having role-switching in our study was the length of the game and its cognitive demands, which we feared would lead to player exhaustion if roles were switched mid-game. In future iterations, we aim to improve the design by allowing players to save their progress and switch roles during subsequent sessions. This could offer a more immersive experience, as players would gain perspectives from both the misinformation creator and debunker roles. Additionally, we aim to introduce new modes, such as a family mode, where players interact with two LLM-simulated characters. This option would reduce cognitive load while maintaining engagement.
@@
The current game is also limited in its sole focus on a pandemic scenario. In reality, misinformation spans multiple domains, with health-related misinformation able to influence political events such as elections. Our game only addressed text-based misinformation, while visual and video-based misinformation pose even greater challenges and are harder to detect. Future work could include multimedia content, such as images and videos, to more accurately simulate the diverse forms of misinformation that exist in the real world.
@@
While LLMs like ChatGPT-4o were used to simulate human reactions in the game, these models do not fully replicate the complexities of human behavior. Human reactions are often nuanced and influenced by multiple factors, including culture, history, and personal experience. LLM-generated characters may oversimplify human emotions and fail to grasp the full context of certain situations. Despite us using an advanced ChatGPT-4o model, it may have missed these subtleties, leading to interactions that feel artificial or incomplete. For example, in the gameplay, players employ strong emotional manipulation strategies to provoke specific responses. However, the LLM-generated characters do not react as expected, where participants reported feeling frustrated when their emotional manipulation strategies did not yield the anticipated reactions. This can limit the diversity and depth of the simulated interactions and detract from the realism and fairness of the experience for certain audiences.
LLMs are trained on large datasets that may contain intrinsic biases, which can manifest themselves in unintended ways during gameplay(Kasneci et al., 2023). Studies have shown that LLMs are prone to so-called ”hallucinations” and can reflect stereotypes or skewed perspectives, which could affect how certain characters respond in the game(Xie et al., 2024). For instance, if biased training data influence a simulated character’s reaction, it may inadvertently reinforce player stereotypes about certain groups of people.
@@ 7. Conclusion
Game-based approaches have shown great promise as tools for inoculating individuals against the tactics commonly used to spread misinformation. Most existing games in this domain are single-player games which offer players limited, predefined choices. While this design reduces cognitive load, it often results in interactions which feel less natural and engaging. In response, we designed a two-player, PvP game that pits a misinformation creator against a misinformation stopper. By integrating LLM-powered personas to evaluate player outputs and provide real-time feedback, we created a more open-ended and immersive experience. We found that the game we developed effectively improved players’ media literacy. Participants demonstrated an enhanced ability to evaluate and analyze media content, identify unreliable or misleading information, and employ effective counter-misinformation strategies. Moreover, the game’s engaging mechanics, combined with the competitive element, motivated players to learn from both their own strategies and those of their opponents. These findings suggest that integrating dynamic feedback systems and competitive gameplay elements into misinformation education games offers a compelling method to deepen users’ engagement, while also improving their critical media skills. Future research can build on these insights to explore other forms of interactive learning environments, focusing on diverse player experiences and varying misinformation challenges.
@@ Appendix A Appendix
A.1. Demographic Information of Participants
@@
A.2. In-game News
Below, we present the ”News” in four rounds.
Round 1 Dr. Amy claims that the herbal ingredients in R medicine have antiviral, anti-inflammatory, and immune-regulating properties that could potentially treat the Zinc virus. Dr. Amy has a medical background from a top university, has worked at the largest hospital in Southland City for 20 years, and is a member of the Academy of Medicine in South Country. She recently joined a newly established traditional medicine research institution and has started advocating for traditional medicine. She has a large online following, and many people trust her words. Her findings have gained support from those worried about the virus. Eternal Wellness claims that 20% of R medicine sales will go toward Zinc virus research and prevention, which has boosted public goodwill. The product, being affordable, has sold out in many places. This newspaper, known for its anti-corruption stance, has further enhanced Dr. Amy’s positive image.
Round 2 More patients are dying in the city, with some cases attributed to severe Zinc virus infections. Among those who have died, a few had used the new medicine Max, raising concerns among the public. One widow shared a story on social media about how her husband contracted the Zinc virus, his time in the hospital, and the treatment he received. She suspects that Max was ineffective and believes it may have caused renal impairment, eventually leading to his death. She claims, “He was given a medication we demanded he NOT receive, and his health quickly went downhill,” ultimately resulting in him being “on a ventilator working most of the time at 100%.” This story has gained significant attention on the largest social media platform, though it remains unclear how widespread these experiences are or whether there is any scientific basis for her claims.
Round 3 A journalist investigating the case has discovered that the traditional medicine institution where Dr. Amy works received significant funding from billionaire Jack. Although this institution is associated with a prestigious research institute, there are concerns that it might be leveraging the institute’s name without a true affiliation. Adding to the complexity, billionaire Jack’s ex-wife owns a company that produces and promotes traditional medicine products. These connections have led to suspicions that the recent promotion of traditional medicine might be part of a strategic collaboration between billionaire Jack and his ex-wife, raising questions about possible financial motivations behind this push.
Round 4 A video circulating on a popular short video platform claims that a doctor practicing alternative medicine and R was targeted and killed to protect the financial interests of ”Big Pharma.” In the video, his friend, an herbalist with over 14,000 followers, states, ”They are cracking down because Big Pharma is losing money.” She adds, ”My friend was healthy and not suicidal, but we need to stand up because this, this is going to get us all killed.” The video has gained significant traction, with 100K views and shared over 30K times. As a result, more people are rallying in favour of traditional medicine, believing it has been effective throughout history and is a proud part of their culture. They advocate for the use of traditional medicine and oppose new drugs, seeing resistance to traditional medicine as an attack on their cultural heritage.

### v2

@@ 6.3.1. Balance Between Freedom and Guidance
In this game, we aimed to go beyond the typical choice-based approach in misinformation education games by enabling free-form input. We found this approach triggers reflection, which helps to build hands-on experience and make the game more enjoyable. Yet, we also found that it relies on players’ existing knowledge of misinformation. For example, players might incorporate unverified information they’ve encountered on social media into the game, which is specifically problematic for the Journalist role. While we provided the players with comprehensible instructions to guide their role’s actions (how to act as a Journalist or an Influencer), it would be better to incorporate more context-specific tips in each stage of the game to help users explore different ways of winning the game and deepening their learning. We suggest using the approach used in the free-input educational interventions (e.g. (Bryfczynski, 2012)) to build clear, understandable criteria for free-form answers. These will not stop creativity but help people tailor their answers to the context of the game. In addition, to help guide players to an understanding of their roles, we could add a preliminary stage to the game. For example, the Journalist role could be introduced to fact-checking and debunking guidelines (EDMO, 2024; Agence France-Presse, 2023; European External Action Service, 2021), with a brief comprehension check before the main game.
@@ 6.3.2. Replayability and Feedback
One of the critical challenges in serious games is maintaining replayability, as this is important to facilitate the learning process (Adetunji and Ade-Ibijola, 2024) and making interventions more sustainable (Silveira, 2016). Moreover, a lack of replayability in educational games can limit both educational and behavioral change (Epstein et al., 2021). We envisioned the designed elements that could encourage multiple playthroughs, such as two distinct player roles and a free-form input mechanism that broader decision space. Future interventions could further enhance replayability by introducing a wider range of characters to represent public opinion. In our game, we observed that players were more engaging when characters resonates with them personally. By introducing more characters, or by allowing players to customize characters to better reflect their own experience, the game could encourage players to return and interact with new characters. The result would be a more engaging experience. For instance, research has shown that debunking misinformation often occurs within families and can cause conflict (Scott et al., 2023). Players could customize a character based on their previous experiences with family members, thereby practicing their own debunking strategies without risk of conflict with family members.
Another way to improve the game’s educational value is feedback. Research has found that debriefing is a crucial opportunity for players to process and integrate their learning experiences(Crookall, 2014; Leder et al., 2024; Barzilai and Stadtler, 2024). After gameplay, we suggest arranging debriefing sessions that allow players to review their strategies, assess their effectiveness, and receive constructive feedback, potentially improving learning outcomes. For instance, after a session focused on combating misinformation, a post-game review might present an ideal debunking response or a well-supported counterargument. Such structured reflection enables players to internalize lessons and increases the likelihood that they will re-enter the game with newly gained insights, thereby reinforcing both learning and replayability.
@@ 6.4. Limitations and Future Work
6.4.1. Limitied Demographics Diversity
The generalizability of our results is constrained by participant demographics. Prior research has established cultural variations in susceptibility and responses to misinformation. For instance, Roozenbeek et al. (Roozenbeek et al., 2020a) indicated that Mexican and Spanish users exhibited higher trust in misinformation than those from Ireland, the UK, and the USA. In contrast, other studies highlight that non-Western populations may be more responsive to misinformation interventions compared to Western counterparts(Noman et al., 2024). Within our East Asian sample, we observed a noticeable resonance with traditional medicine narratives. While traditional and alternative medicines hold varying levels of acceptance globally(Swire-Thompson et al., 2020), this particular cultural context may have facilitated the effectiveness of our intervention when these themes were present. The positive influence of this resonance on player engagement suggests that culturally relevant themes can impact this type of intervention’s engagement and potentially learning outcomes. Therefore, future research should investigate the applicability of these findings in other cultural settings. We recommend tailoring culturally relevant themes for different target populations based on their specific interests and vulnerabilities. For example, given the higher susceptibility of Western audiences to politically aligned misinformation, particularly during election periods (Humprecht et al., 2020), the game could be reframed around political theme. In such a version, the Influencer could be tasked with spreading partisan or misleading claims, while the Journalist attempts to debunk them. Background news can be adapted to high-stakes topics such as taxation, immigration, or public policy, while the game mechanics can remain unchanged. The flexibility of the game design supports broader applicability across domains and audiences while preserving its educational objectives.
@@ 6.4.1. Limitied Demographics Diversity
Previous studies also suggest that certain populations may face greater challenges in being able to critically evaluate information. For example, a large-scale study observed that Asian individuals encounter more difficulties in assessing health information from social media compared to other populations (Chandrasekaran et al., 2024). In addition, they’re more likely to incorporate social media information into their health-related decisions, potentially increasing their susceptibility to misinformation (Chandrasekaran et al., 2024). Thus, cultural background of our participants can potentially make them more susceptible to misinformation than other populations.
@@
Our sample was relatively homogeneous in age. A recent meta-analysis of articles about different intervention approaches showed that neither age nor gender significantly impacts the effectiveness of media literacy interventions (Lu et al., 2024). However, previous work has suggested that media literacy interventions designed for certain age groups (e.g., older adults and adolescents) achieved greater effects(Moore and Hancock, 2022; Hartwig et al., 2024). Future work should determine if our approach is efficient in other age groups and, if necessary, tailor scenarios to suit the different age groups.
6.4.2. Study Design
@@
Our study provided only a one-time intervention and observed immediate learning effects; previous work showed that even a one-time interaction with an educational game can provide long-term improvement in misinformation recognition. For instance, Maertens et al. tested the game “Bad News” and found that inoculation effects lasted for at least 13 weeks. This suggests the potential for the long-term effectiveness of active inoculation interventions with regular assessment(Maertens et al., 2021). Still, future research should include multiple time points to assess the long-term effectiveness of our game intervention. There should also be comparisons between one-time and multiple play sessions, with explorations of the impacts of players assuming different roles within the games.
6.4.3. Game Design
@@
The current game has a limited focus on a pandemic scenario. In reality, misinformation spans multiple domains, with health misinformation able to influence political events such as elections; therefore, future works should focus on incorporating scenarios reflecting the connection between different domains of misinformation; it should also incorporate participatory codesign of the topics with potential participants to be sure the senarious reflects a real use-cases of misinformation encounter (Cook et al., 2023). Our game only addressed text-based misinformation, while visual and video-based misinformation pose even greater challenges and are harder to detect. Future work could include multimedia content, such as images and videos, to more accurately simulate the diverse forms of misinformation that exist in the real world.
In this study, each player was limited to a single role, either an Influencer or a Journalist. This resulted in different learning experiences depending on their assigned role. The primary reason for not having role-switching in our study was the length of the game and its cognitive demands, which we feared would lead to player exhaustion if roles were switched mid-game. In future iterations, we aim to improve the design by allowing players to save their progress and switch roles during subsequent sessions. This could offer a more immersive experience, as players would gain perspectives from both the Influencer and Journalist roles.
@@
6.4.4. LLM Biases and Hallucination Risks
Although our game is set in a fictional country with fictional personas, we did not clearly define their sociocultural backgrounds, which may introduce biases in the representation of these personas. Prior research indicates that LLMs often reflect the cultural norms dominant in their training data, which are largely based on English-language materials from Western contexts (Liu, 2024; Salminen et al., 2024). In addition, OpenAI specifically fine-tunes its models to avoid providing misinformation and to give answers to users from the perspective of scientific consensus(OpenAI Achiam et al., 2023). This inherent social value alignment and moderation can introduce potential positive bias towards westernised, pro-scientific arguments. For example, these personas may disproportionately favor pro-science, anti-conspiracy, and neutral-toned arguments, even when such arguments are not strongly supported by evidence in the game. This tendency undermines the realism of the misinformation spreading and correction processes depicted in the game and may unintentionally encourage players to prioritize persuasive tactics emphasizing surface features (e.g., fluency, neutrality, scientific-sounding language) over substantive content (accuracy, logical coherence).
Furthermore, LLMs responses are prone to hallucinations, which can undermine the game’s educational effectiveness(Xie et al., 2024; Park et al., 2023). For instance, if a persona fabricates reasons to trust a player’s arguments, players may incorrectly infer that their argumentation techniques are effective, inadvertently internalizing flawed reasoning strategies. Over time, this may negatively impact the game’s intended learning outcomes. Additionally, hallucinations may diminish players’ trust in the learning process. Players may struggle to understand why their actions are being praised or criticised when personas provide verbose, ambiguous, or logically inconsistent feedback. As shown by Kaate et al.(Kaate et al., 2025), users are often frustrated with long, unclear, or irrelevant AI-generated responses. Thus, it could negatively affect both learning outcomes and motivation. To mitigate these issues, future work should consider explicitly defining the sociocultural backgrounds and value orientations of LLM personas to reduce biases. We also recommend adding features like uncertainty cues or clarification prompts to acknowledge system limitations. Such designs can help players critically evaluate feedback and recognize potential inaccuracies (Kaate et al., 2025).
It is also important to recognize that LLMs do not fully replicate the complexities of human behavior. Human reactions are influenced by multiple factors, including culture, history, and personal experience. AI-driven characters may oversimplify human emotions and fail to grasp the full context of certain situations. For example, during gameplay, we observed that players employed strong emotional manipulation strategies intended to provoke specific responses, but the characters did not react as anticipated. Participants reported frustration when their strategies failed to yield expected reactions, which reflects the discrepancy between the simulated interactions and realistic human reactions. To minimize these inconsistencies, we implemented strict prompt engineering protocols to define the AI characters’ output parameters (See details in 3.3). Our process also involved multiple internal testing iterations and two pilot gameplay sessions, from which we analyzed LLMs output and iteratively refined prompts for enhanced consistency. Despite these measures, occasional variability in LLM-generated responses persisted, potentially affecting participant engagement and the learning outcomes. As prior research suggests that creating complex game rules and mechanics with LLMs requires extensive fine-tuning and human intervention (Zhang et al., 2025). Future work could focus on developing robust prompt engineering practices, clear guidelines for human intervention, and standardized methodologies for monitoring and evaluating LLM performance.
Lastly, with novel technology, the LLMs presents potential risks, such as inconsistent responses, reinforcement of stereotypes, limited cultural representation, and hallucinations (Salminen et al., 2024). In our study, LLMs were used as in-game characters to simulate public opinion and provide feedback to players rather than correcting misinformation. This framing helped reduce some risks, as the personas were not positioned as truth arbiters. Furthermore, the game was designed with multiple elements, including background news content, PvP mechanics, and free-form input that contributed to learning outcomes beyond the LLMs feedback. As our primary focus was on player interactions and strategy development, we did not systematically assess the risks posed by LLM-powered feedback. Nevertheless, given that these characters influence players’ perceptions, reasoning, and motivation, we encourage future work to examine the educational impact, limitations, and potential harms of LLM-powered feedback more rigorously in similar game-based learning environments.
@@ 7. Conclusion
Game-based approaches have shown great promise as tools for inoculating individuals against the tactics commonly used to spread misinformation. Most existing games in this domain are single-player games which offer players limited, predefined choices. While this design reduces cognitive load, it often results in interactions which feel less natural and engaging. In response, we designed a two-player, PvP game that pits a misinformation creator against a misinformation stopper. By integrating LLM-powered characters to evaluate player outputs and provide real-time feedback, we created a more open-ended and immersive experience. We found that the game we developed effectively improved players’ media literacy. Participants demonstrated an enhanced ability to evaluate and analyze media content, identify unreliable or misleading information, and employ effective counter-misinformation strategies. Moreover, the game’s engaging mechanics, combined with the competitive element, motivated players to learn from both their own strategies and those of their opponents. These findings suggest that integrating dynamic feedback systems and competitive gameplay elements into misinformation education games offers a compelling method to deepen users’ engagement, while also improving their critical media skills. Future research can build on these insights to explore other forms of interactive learning environments, focusing on diverse player experiences and varying misinformation challenges.
Acknowledgments
Thanks to Jiaming Zhou for support with conceptualization. This work was supported by the TDG Teaching Development Grant (Proj 6000901), the TRS Theme-based Research Scheme (T45-205/21-N), and the Luxembourg National Research Fund (REMEDIS, REgulatory and other solutions to MitigatE online DISinformation (INTER/FNRS/21/16554939)).
@@ Appendix A Appendix
A.1. Demographic Information of Participants in Game Intervention
@@
A.2. Control Group Demographic Information of Participants
Table 4. Control group’s demographic details of participants (N=50)
Number Gender Age Education Level Profession
N1 Male 24 Master’s degree Design
N2 Female 26 Master’s degree Business Analysis
N3 Female 28 Master’s degree Design
N4 Female 23 Bachelor’s degree Chemistry
N5 Female 32 Bachelor’s degree Business
N6 Male 24 Master’s degree Computer science
N7 Male 23 Master’s degree Design
N8 Male 29 Master’s degree IT
N9 Female 24 Bachelor’s degree Bioengineering
N10 Female 21 Bachelor’s degree Auditing
N11 Female 21 Bachelor’s degree Animal biology
N12 Male 21 Bachelor’s degree Mechanical engineering
N13 Female 32 Bachelor’s degree Finance
N14 Female 26 Bachelor’s degree Digital Media Arts
N15 Female 24 Master’s degree Law
N16 Female 26 Master’s degree English
N17 Female 38 Bachelor’s degree Accounting
N18 Female 25 Master’s degree Biomedical science
N19 Female 41 Master’s degree Clinical medicine
N20 Male 26 Bachelor’s degree IT
N21 Female 23 Bachelor’s degree Journalism
N22 Female 23 Bachelor’s degree Medical science
N23 Male 24 Master’s degree Architecture
N24 Female 24 Bachelor’s degree Barista
N25 Female 22 Master’s degree Law
N26 Female 25 Master’s degree Media
N27 Male 22 Bachelor’s degree Chemistry
N28 Male 34 Bachelor’s degree Intelligent manufacturing engineering technology
N29 Female 23 Master’s degree Sociology
N30 Female 21 Bachelor’s degree Computer Science and Technology
N31 Female 22 Bachelor’s degree Materials science
N32 Male 22 Bachelor’s degree Rehabilitation Therapeutics
N33 Male 42 Bachelor’s degree Mechanical engineering
N34 Female 28 Bachelor’s degree Finance
N35 Female 25 Master’s degree Computational design
N36 Female 38 Associate degree Accounting
N37 Male 36 Bachelor’s degree Mechanical engineering
N38 Female 26 Master’s degree Data science
N39 Male 36 Associate degree Mechanical engineering
N40 Male 28 Associate degree Telecommunications engineering
N41 Female 21 Bachelor’s degree Financial management
N42 Female 39 Bachelor’s degree Electrical Engineering and Automation
N43 Female 35 Associate degree Chinese language and literature
N44 Female 25 Associate degree Pre-primary education
N45 Female 42 Bachelor’s degree Visual communication design
N46 Female 24 Bachelor’s degree Primary education
N47 Female 21 Master’s degree Design
N48 Male 40 Bachelor’s degree Art
N49 Male 21 Bachelor’s degree Intelligent Manufacturing Engineering Technology
N50 Male 27 Associate degree Medical science
Table 4. Demographic details of Participants (continued)
A.3. Prompt Design
Story Context: The Southland, with a 6,000-year history, boasts rich natural resources, diverse ecosystems, and a culture that values liberty, free expression, and media independence. Historically, Southland has been known for its renowned doctors and pharmacists, and its people are proud of their traditional medicine, which differs from modern methods and has recently been debated. Despite its history of dealing with epidemics, Southland was caught off guard by the mysterious Zinc Virus. Believed to stem from ecological imbalances in southern rainforests, the virus causes high fever, respiratory issues, and immune collapse. Its high transmissibility led to a swift outbreak in the capital, Southport, triggering a public health crisis. The scientific community cannot yet provide confident and effective treatments due to the uncertainty surrounding the virus. Hospitals offer only standard treatments. In the absence of definitive medical solutions, people turn to existing medications for similar symptoms. The outbreak led to deserted streets, closed public spaces, and a reliance on social media for information, which also spread rumors and misinformation. Scientists raced to find treatments while traditional medicine’s popularity surges despite scientific doubts about its effectiveness.
Role Description:Your role is to simulate how the five personas react to each piece of information they receive. The game will be divided into four rounds. In each round, you already knew about the background events happening I gave you in 1.3. In the first round, you will first receive a piece of misinformation, and you need to simulate public opinion and provide a response. Then, you will receive a piece of information that counters the previous misinformation, and you will need to simulate public opinion’s reaction to this new information as well. After that, the first round ends and moves to the second round, which follows the same process as the first: you will first receive a new piece of misinformation, simulate public opinion and provide a response, then receive a counter-misinformation piece and simulate the personas’ reactions to it, again providing a public opinion response.
Persona 1: Emily. 16. Female. High School Student. Political party affiliation: No. Extraversion: High - Outgoing and talkative, enjoys socializing with friends. Agreeableness: High - Trusting and kind, easily believe what friends and influencers share. Conscientiousness: Low - Less responsible and thorough, tends to be impulsive and carefree. Neuroticism: Medium - Sometimes anxious about fitting in and being accepted by peers. Openness: Low - Limited exposure to diverse experiences, prefers familiar and popular content. Relies heavily on social media for information.Has limited critical thinking skills and media literacy. Tends to believe information shared by friends and influencers without verifying facts. Limited exposure to diverse sources of information.
Persona 2: Alex. 36. Male. Project Manager in a Corporate Firm. Undergraduate. Political party affiliation: Strongly support Liberal. Extraversion: Medium - Enjoys social activities but also values alone time for work and personal projects. Agreeableness: Medium - Generally trusting and cooperative, but can be skeptical of new information. Conscientiousness: Medium - Balances responsibilities but can be hasty in decision-making. Neuroticism: Low - Generally calm and composed, rarely anxious or stressed. Openness: Medium - Open to new experiences but sometimes prefers convenience over exploration. Busy with work responsibilities and managing projects, often skims through news during short breaks. Follows news via quick-read apps and social media. Often shares articles based on headlines without reading fully. Some critical thinking skills but lacks depth in media literacy.
Persona 3: Maria. 46. Female. Housewife. High school Degree. Political party affiliation:No. Extraversion: High - Very outgoing and enjoys talking with different people and relatives. Agreeableness: High - Generous and kind. Conscientiousness: High - Very responsible and thorough, values their culture and proud to be Southland’s citizens and long history. Neuroticism: High - Highly anxious, extremely anxious when encounter pressure. Openness: low - support traditional and conservative values. Regularly reads a variety of news sources, both local and international, loves to learn knowledge about health and wellness, especially natural remedies. Critical thinker with a low understanding of media bias. Not susceptible to misinformation and nearly blindly believe natural remedies, afraid of hospital.
Persona 4: John. 78. Male. Retired Nurse. Master’s Degree. Political party affiliation: moderate support conservative. Extraversion: Low - Prefers quiet activities and smaller gatherings over large social events. Agreeableness: Medium - Generally kind and cooperative but maintains a healthy level of skepticism. Conscientiousness: High - Very responsible and thorough, values consistency. Neuroticism: Medium - Occasionally anxious about new technologies and online trends. Openness: Medium - Open to new ideas but prefers well-established and trusted sources. Relies on traditional media (TV, newspapers) but is starting to use social media. Skeptical of new technologies and online information. Prefers information from established sources but may not be up-to-date with digital literacy. Have knowledge of health information because of occupation.
Persona 5: Sophia. 27. Female. Political party affiliation: Strongly support Conservative. Freelance Graphic Designer. High School. Extraversion: High - Very outgoing and active on social media, enjoys engaging with others. Agreeableness: High - Trusting and kind, values community and cooperation. Conscientiousness: Medium - Generally responsible but can be impulsive online. Neuroticism: Medium - Sometimes anxious about online perceptions and fitting in. Openness: High - Intellectual and artistic, enjoys exploring new ideas and content. Highly active on social media and tech-savvy. Frequently encounters various information online, including memes and viral content. Has basic media literacy skills but sometimes struggles to differentiate satire from misinformation. Interested in learning more about media literacy and fact-checking.
Response Format: here is the format for each one.
  • Persona 1 Basic Details (e.g., age, gender, occupation) Persona Reaction: Trust Level Score:
  • Persona 2 Basic Details (e.g., age, gender, occupation) Persona Reaction: Trust Level Score:
  • Persona 3 Basic Details (e.g., age, gender, occupation) Persona Reaction: Trust Level Score:
  • Persona 4 Basic Details (e.g., age, gender, occupation) Persona Reaction: Trust Level Score:
  • Persona 5 Basic Details (e.g., age, gender, occupation) Persona Reaction: Trust Level Score:
Trust Level Score Criteria:
  • Score 0-2: Very low trust, these personas are highly unlikely to believe the fake news and will typically seek verification.
  • Score 3-4: Low trust, these personas are cautious and tend to verify information before believing it.
  • Score 5-6: Moderate trust, these personas may sometimes be misled, depending on the source and context.
  • Score 7-8: High trust, these personas are easily misled and usually do not verify the information.
  • Score 9-10: Very high trust, these personas almost unquestioningly believe and spread fake news.
In each round, you should consider the event and human players’ output to evaluate. Consider more like a human evaluator, giving more diversity score distribution.
A.4. Game Mechanics Design
Scoring System: In each round, each LLM character evaluates the messages provided by the players and provides an opinion and trust level score on a 10-point scale:
  • 10: The character fully trusts the misinformation published by the Influencer.
  • 0: The character completely trusts the debunking response published by the Journalist.
  • If the final score is above 5, Influencer wins.
  • If the final score is below 5, Journalist wins.